# Language Models Optimized to Fool Detectors Still Have a Distinct Style (And How to Change It)

### v1

@@ 6 Experiments
The goal of our main experimental evaluations is to: (1) demonstrate that our system can evade machine-text detectors, even in multi-sample settings; (2) estimate the upper bound of distributional overlap using AURA; and (3) show that our paraphraser outperforms existing style-aware methods.
Machine-Text Detection as the Samples Size Grows
In this section, we study whether machine-text detectors are robust against various attacks as the sample size grows. Although two distributions may appear indistinguishable on a per-sample basis, their differences become more apparent as the number of samples increases. For each detector, we compute the score $s_{i}$ by taking the sample mean of its outputs over $n$ samples. For each value of $n$, we report the best score achieved across the detectors described in §missing 5 for a pessimistic estimate of the detectability of each attack. These results are shown in Figure 2. We find that our approach is the least detectable, even in domains for which it was not trained (Amazon and Reddit). Although our approach transfers well to Amazon, we find that it becomes detectable with just $5$ samples in the Blogs domain. We attribute this to the large domain mismatch between the training data (Reddit), favoring informal social media text, and the more structured, formal Blogs text. To better understand the differences between each detector, we break down the per-detector performance for our method and Mistral-7B-DPO-FastDetectGPT, the next least detectable approach, on Reddit in Figure 3. The results highlight that although Mistral-7B-DPO-FastDetectGPT is robust against FastDetectGPT, the detector it was explicitly optimized against, as well as others that rely on similar token-level features, it remains easily identifiable by StyleDetect, which leverages writing style. In contrast, our approach is more universally undetectable across all detectors tested although not being explicitly optimized against any of them. Finally, in Table 3, we show the semantic similarity and the character edit distance of each approach. We find that our approach preserves the meaning of the original text (similarity $>0.85$), while making on average $+43$ more character edits than regular paraphrasing. We attribute this increase in edits to the necessary constraint of following the target author’s writing style.
Methods $\rightarrow$ Prompting Paraphrasing TinyStyler Ours
Edit Distance 134.0 156.6 212.6 199.1
Semantic Sim. 0.90 0.93 0.78 0.86
Table 3: Character edit distance, and semantic similarity of the different methods evaluated. Results averaged across datasets, for full breakdown see Appendix B.
@@
Figure 2: Detection performance (AUROC(10)) of the strongest detector for each sample size and method combination. Our method is the least detectable across all three domains, including Amazon and Blogs, which were not seen during training.
@@
Figure 3: Performance of various detectors as the sample size increases (left: Mistral-7B-DPO-FastDetectGPT, right: Ours). Our approach is consistently harder to detect across all detectors. Mistral-7B-DPO-FastDetectGPT becomes detectable with just 5 samples, while our approach remains robust up to 50. We report the performance of all detectors, evaluated on all methods and all datasets in Appendix A.
Estimating the Distributional Overlap
Method $\downarrow$ / AURA $\rightarrow$ Classifier $\uparrow$ Token Dist. $\uparrow$ Style $\uparrow$ Average
Baseline 0.63 0.08 0.00 0.24
Mistral-7B-DPO-FastDetectGPT 0.67 0.01 0.00 0.23
Paraphrasing 0.62 0.26 0.00 0.29
Prompting 0.66 0.09 0.00 0.25
TinyStyler 0.01 0.00 0.00 0.00
Ours 0.78 0.27 0.78 0.61
Table 4: AURA calculated across three classes of detectors: classifiers (OpenAI and RADAR), those that use the predicted distributions (Rank, LogRank, FastDetectGPT and Binoculars), and those that use style (StyleDetect). Our approach is the best closes the gap between the human and machine distributions according to all detector classes.
In the previous section, we showed that while paraphrasing and Mistral-7B-DPO-FastDetectGPT degrade detector performance in the single-sample setting, this effect diminishes as more samples become available. This raises the question: how close are the underlying distributions? To answer this, we use AURA (§missing 4) to quantify the distributional overlap. We divide the detectors into three categories: those that train a classifier (OpenAI and RADAR), those that use the predicted token distributions (Rank, LogRank, FastDetectGPT and Binoculars), and those that use features indicative of writing style (StyleDetect). For each $n\in\{1,5,10,25,50,100,200\}$, we calculate the AUROC achieved by the detectors on the Reddit dataset, and take the maximum AUROC across each category to be the representative value for that category. Using these curves, we calculate the AURA metric across the three categories and report our findings in Table 4. Higher AURA scores indicate greater similarity between the detector score distributions of human- and machine-generated text, suggesting that the detector finds the two distributions harder to distinguish. We find that while most approaches (except TinyStyler) appear similar to trained classifiers, every approach except ours remain distinguishable according to its writing style. Surprisingly, we find that simple paraphrasing is the second-most robust attack, outperforming even Mistral-7B-DPO-FastDetectGPT, a finding that aligns with prior work (Krishna et al., 2023; Sadasivan et al., 2025). We find that although Mistral-7B-DPO-FastDetectGPT is capable of avoiding FastDetectGPT and Binoculars, it fails to evade Rank and LogRank (Figure 3), suggesting that detectors that use the predicted distributions may rely on distinct, non-overlapping features. Finally, our approach is weakest in avoiding detection by methods that rely on the predicted distributions.
Style-Aware Paraphrasing Performance
@@
Figure 5: Performance of the best detector for each sample size evaluated on outputs of our system with, and without DPO. DPO helps maintain the generations undetectable.
In this section, we compare the performance of our style-aware paraphraser to TinyStyler, a recent method for author-conditioned style transfer. We evaluate both systems on the Reddit dataset described in §missing 5.1. We find that our approach improves upon the stylistic similarity achieved by TinyStyler by $+0.12$ (from $0.71$ to $0.83$), and the semantic similarity by $0.09$ (from $0.74$ to $0.83$).
7 Ablations
@@ 8 Related Works
author=Rafael,color=green!40,size=,inline,]I will make this better.
@@
Another line of work aims to defeat machine-text detectors, either through paraphrasing (Krishna et al., 2023; Sadasivan et al., 2025), by prompt optimization (Lu et al., 2024), by adding a single space in the generation (Cai and Cui, 2023), with homoglyphs (Creo and Pudasaini, 2025), or more recently by post-training LLMs with DPO to prefer generations that evade detection (Nicks et al., 2024; Wang et al., 2025). However, we show that these approaches fail to close the gap between human and machine-text distributions, as they primarily manipulate surface-level features without altering the underlying writing style (§missing 2). In contrast, our method is the first to jointly optimize for author-specific human writing styles and against the surface-level features exploited by most detectors.
@@ Appendix B Breakdown of Edit Distance and Semantic Similarity by Dataset
Methods $\rightarrow$ Prompting Paraphrasing TinyStyler Ours
@@
Edit Distance 107.33 122.74 158.78 169.57
Semantic Sim. 0.87 0.90 0.77 0.82
@@
Edit Distance 128.06 143.12 209.61 178.01
Semantic Sim. 0.94 0.96 0.84 0.90
@@
Edit Distance 166.75 203.85 269.35 249.68
Semantic Sim. 0.90 0.92 0.73 0.85
@@ Appendix C Training Hyperparameters and Compute Resources
Our system is parametrized using Mistral-7B, trained for $1$ epoch on the Reddit dataset described in §missing 5.1 with a constant learning rate of $2e^{-5}$, using LoRA (Hu et al., 2021) for efficient fine-tuning, setting $r=32$, $\alpha=64$, and $d=0.1$. For the preference-tuning stage, we train our system with $\beta=5$, and a constant learning rate of $1e^{-6}$. For Mistral-7B-DPO-FastDetectGPT, we set $\beta=0.1$.
@@ Appendix D Dataset details
We train our system on the Reddit Million Users Dataset, which contains comments from 1 million authors (Khan et al., 2021). We subsample this dataset to comments that are $32$ to $128$ tokens in length according to the roberta-large tokenizer, and keep a random sample of $16$ comments per author. To ensure that the authors are stylistically diverse while meeting our computational constraints, we further subsample the dataset using stratified sampling in stylistic space. Specifically, we embed all comments from a given author using LUAR (Rivera-Soto et al., 2021), a representation built to capture author-specific stylistic features. We then apply Affinity Propagation (Frey and Dueck, 2007) to cluster the authors, sampling evenly across clusters until reaching $63{,}184$ authors which was computationally tractable given our resources. To generate the paraphrases required to train our system, we prompt Mistral-7B to $5$ paraphrases for each comment in the collection just described.
@@
We evaluate our approach across three domains: Reddit, Amazon, and Blogs. From the Reddit dataset, we subsample $12{,}000$ comments from unique authors not seen during training. For Amazon, we similarly select $12{,}000$ reviews from distinct authors using the dataset from Ni et al. (2019). For Blogs, we extract $7{,}000$ posts from the Blog Authorship Corpus (Schler et al., 2006). To generate machine text, we prompt one of Mistral-7B, gpt-4o-mini, or Llama-3, chosen uniformly at random, to create new comments, reviews, or blog snippets (see prompts in Appendix F). Each baseline described in §missing 5.2 is then applied to modify this generated text to evade detection. The only exception is Mistral-7B-DPO-FastDetectGPT, which generates the text directly, rather than modifying pre-existing outputs. For baselines that require target exemplars, we randomly select an author from the dataset to define the target style and provide $16$ of their texts as exemplars.
@@ F.2 Paraphrasing with GPT-4
For the GPT-4 paraphrasing baseline described in §missing 5.2, we use the following prompt:
@@ F.3 Generating Machine-text
To generate the machine-text samples for the machine-text detection evaluation dataset described in §missing 5.1, we prompt one of Mistral-7B, Phi-3, or Llama-3, uniformly at random, to generate responses to Reddit comments, new Amazon reviews, or new Blog snippets. We found that specifying the number of words in the prompt better contorlled the length of the generations.
@@ F.4 Style-paraphrasing Prompt
The following is the main prompt we use to instruction-tune our system, and for the GPT-4 paraphrasing baseline described in §missing 5.2:

### v2

@@ 5 Experiments
The goal of our main experimental evaluations is to: (1) demonstrate that our system can best evades machine-text detectors § 5.1; show that our approach best closes the gap between human-written and machine-generated styles § 5.2 and (3) show that our paraphraser outperforms existing style-aware methods § 5.3.
@@
Figure 2: Detection performance (AUROC(10), lower is better) of the strongest detector for each sample size and method combination. Our detector evasion approach is the least detectable across all three domains, including Amazon and Blogs, which were not seen during training.
@@
Figure 3: Detection performance (AUROC(10), lower is better) of various detectors as the sample size increases (left: Mistral-7B-DPO-FastDetectGPT, right: Ours). Our detector evasion approach is consistently harder to detect across all detectors. Mistral-7B-DPO-FastDetectGPT becomes detectable with just 5 samples, while our approach remains robust up to 50. We report the performance of all detectors, evaluated on all methods and all datasets in Appendix A.
5.1 Machine-Text Detection as the Samples Size Grows
In this section, we study whether machine-text detectors are robust against various attacks as the sample size grows. Although two distributions may appear indistinguishable on a per-sample basis, their differences become more apparent as the number of samples increases. For each detector, we compute the score $s_{i}$ by taking the sample mean of its outputs over $n$ samples. For each value of $n$, we report the best score achieved across the detectors described in § 4 for a pessimistic estimate of the detectability of each attack. These results are shown in Figure 2. We find that our approach is the least detectable, even in domains for which it was not trained (Amazon and Reddit). Although our approach transfers well to Amazon, we find that it becomes detectable with just $5$ samples in the Blogs domain. We attribute this to the large domain mismatch between the training data (Reddit), favoring informal social media text, and the more structured, formal blogs text. To better understand the differences between each detector, we break down the per-detector performance for our method and Mistral-7B-DPO-FastDetectGPT on Reddit in Figure 3. The results highlight that although Mistral-7B-DPO-FastDetectGPT is robust against FastDetectGPT, the detector it was explicitly optimized against, as well as others that rely on similar token-level features, it remains easily identifiable by StyleDetect, which leverages writing style. In contrast, our approach is more universally undetectable across all detectors tested although not being explicitly optimized against any of them. Finally, in Table 3, we show the semantic similarity and the character edit distance of each approach that relies on transforming text. We find that our approach preserves the meaning of the original text (similarity >0.85>0.85), while making on average +43+43 more character edits than regular paraphrasing. We attribute this increase in edits to the necessary constraint of following the target author’s writing style.
Methods $\rightarrow$ Prompting Paraphrasing DIPPER TinyStyler Ours
Edit Distance 134.0 156.6 227.40 212.6 199.1
Semantic Sim. 0.90 0.93 0.84 0.78 0.86
Table 3: Character edit distance, and semantic similarity of the methods that transform text. Results averaged across datasets, for full breakdown see Appendix B.
Figure 4: UMAP (McInnes et al., 2020) projections of representations that capture writing style for comments in the Reddit domain, using LUAR (Rivera-Soto et al., 2021), CISR (Wegmann et al., 2022), and StyleDistance (Patel et al., 2025). Each point corresponds to a document of at most 128 tokens. Our style aware paraphraser better closes the gap between human-written and machine-generated text (compare with ).
5.2 Visualizing the Space of Writing Styles
We now turn to evaluating whether the approaches considered successfully close the gap between the distributions of human-written and machine-generated writing styles. We choose $100$ samples from Reddit generated by each of Mistral-7B-DPO-FastDetectGPT, DIPPER, OUTFOX, and our style-aware-paraphraser at random. This choice of methods covers the main modalities of detection evasion systems, namely, optimization using DPO, prompting, and paraphrasing. We then embed these generations across three different neural representations of writing-style: LUAR (Rivera-Soto et al., 2021), CISR (Wegmann et al., 2022), and StyleDistance (Patel et al., 2025). We show the results of this in Figure 4. We observe that across all three representations of writing style, our method is qualitatively the one that best closes the gap, further reinforcing that optimization using DPO, prompting, and paraphrasing are insufficient.
5.3 Style-Aware Paraphrasing Performance
@@
Figure 6: Performance of the best detector on Reddit for each sample size evaluated on outputs of our style aware paraphraser with, and without DPO. DPO helps maintain the generations undetectable.
In this section, we compare the performance of our style-aware paraphraser to TinyStyler, a recent method for author-conditioned style transfer. We evaluate both systems on the Reddit dataset described in § 4.1. We find that our approach improves upon the stylistic similarity achieved by TinyStyler by +0.12+0.12 (from $0.71$ to $0.83$), and the semantic similarity by $0.09$ (from $0.74$ to $0.83$).
5.4 Ablations
@@ 6 Related Works
Another line of work aims to defeat machine-text detectors, either through paraphrasing (Krishna et al., 2023; Sadasivan et al., 2025), by prompt optimization (Lu et al., 2024), by adding a single space in the generation (Cai and Cui, 2023), with homoglyphs (Creo and Pudasaini, 2025), or more recently by post-training LLMs with DPO to prefer generations that evade detection (Nicks et al., 2024; Wang et al., 2025). However, we show that these approaches fail to close the gap between human and machine-text distributions, as they primarily manipulate surface-level features without altering the underlying writing style (§ 2). In contrast, our method is the first to jointly optimize for author-specific human writing styles and against the surface-level features exploited by most detectors.
@@ 7 Conclusion
Why is style a robust feature space?
To give the readers some intuitions of why style might be a robust feature space resistant to prompting and optimization via DPO, we note that the representations used by StyleDistance are trained to identify features indicative of individual low-resource authors. While LLMs might be able to replicate the style of high-resource authors such as Shakespeare, or coarse-grained style categories like formal tone or informal tone, it is difficult for them to generate text in the style of a specific low-resource author whose style might be underrepresented in the training data (long-tails of the distribution).
@@
Reproducibility Statement
The datasets, method implementations, model checkpoints, and experimental scripts, will be released along with the paper: https://anonymous.4open.science/status/style-aware-paraphrasing-BD8E
Ethics Statement
The ability to generate convincing machine-generated text poses a significant risk of abuse. This paper contributes an improved understanding of methods to detect machine-generated text, as well as attacks which may hamper the detection of machine-generated text. By studying such attacks, we contribute a better understanding of the limitations of current state-of-the-art defenses, as well as opening the door to future improvements in machine-text detection techniques. Overall, our findings underscore limitations of previous detection regimes, and at the same time suggest that certain feature spaces may be inherently more robust for detection.
@@ Appendix B Breakdown of Edit Distance and Semantic Similarity by Dataset
Methods $\rightarrow$ Prompting Paraphrasing DIPPER TinyStyler Ours
@@
Edit Distance 107.33 122.74 168.02 158.78 169.57
Semantic Sim. 0.87 0.90 0.76 0.77 0.82
@@
Edit Distance 128.06 143.12 223.55 209.61 178.01
Semantic Sim. 0.94 0.96 0.96 0.84 0.90
@@
Edit Distance 166.75 203.85 290.62 269.35 249.68
Semantic Sim. 0.90 0.92 0.81 0.73 0.85
@@ Appendix C Training Hyperparameters and Compute Resources
Our system is parametrized using Mistral-7B, trained for $1$ epoch on the Reddit dataset described in § 4.1 with a constant learning rate of $2e^{-5}$, using LoRA (Hu et al., 2021) for efficient fine-tuning, setting $r=32$, $\alpha=64$, and $d=0.1$. For the preference-tuning stage, we train our system with $\beta=5$, and a constant learning rate of $1e^{-6}$. For Mistral-7B-DPO-FastDetectGPT, we set $\beta=0.1$.
@@ Appendix D Dataset details
We train our system on the Reddit Million Users Dataset, which contains comments from 1 million authors (Khan et al., 2021). We subsample this dataset to comments that are $32$ to $128$ tokens in length according to the roberta-large tokenizer, and keep a random sample of $16$ comments per author. To ensure that the authors are stylistically diverse while meeting our computational constraints, we further subsample the dataset using stratified sampling in stylistic space. Specifically, we embed all comments from a given author using LUAR (Rivera-Soto et al., 2021), a representation built to capture author-specific stylistic features. We then apply Affinity Propagation (Frey and Dueck, 2007) to cluster the authors, sampling evenly across clusters until reaching 63,18463{,}184 authors which was computationally tractable given our resources. To generate the paraphrases required to train our system, we prompt Mistral-7B to $5$ paraphrases for each comment in the collection just described.
@@
We evaluate our approach across three domains: Reddit, Amazon, and Blogs. From the Reddit dataset, we subsample 12,00012{,}000 comments from unique authors not seen during training. For Amazon, we similarly select 12,00012{,}000 reviews from distinct authors using the dataset from Ni et al. (2019). For Blogs, we extract 7,0007{,}000 posts from the Blog Authorship Corpus (Schler et al., 2006). To generate machine text, we prompt one of Mistral-7B, gpt-4o-mini, or Llama-3, chosen uniformly at random, to create new comments, reviews, or blog snippets (see prompts in Appendix F). Each baseline described in § 4.2 is then applied to modify this generated text to evade detection. The only exception is Mistral-7B-DPO-FastDetectGPT, which generates the text directly, rather than modifying pre-existing outputs. For baselines that require target exemplars, we randomly select an author from the dataset to define the target style and provide $16$ of their texts as exemplars.
@@ F.2 Paraphrasing with GPT-4
For the GPT-4 paraphrasing baseline described in § 4.2, we use the following prompt:
@@ F.3 Generating Machine-text
To generate the machine-text samples for the machine-text detection evaluation dataset described in § 4.1, we prompt one of Mistral-7B, Phi-3, or Llama-3, uniformly at random, to generate responses to Reddit comments, new Amazon reviews, or new Blog snippets. We found that specifying the number of words in the prompt better contorlled the length of the generations.
@@ F.4 Style-paraphrasing Prompt
The following is the main prompt we use to instruction-tune our system, and for the GPT-4 paraphrasing baseline described in § 4.2: